Michael is 3 times as old as William and is also 14 years older than William.
How old is William?

Answer by ankor@dixie-net.com(22740)

Michael is 3 times as old as William
m = 3w
:
and is also 14 years older than William.
m = w + 14
:
How old is William?
Repalce m with 3w in the above
3w = w + 14
3w - w = 14
2w = 14
w = 14/2
w = 7 yr is W's age